OpenStudy (anonymous):

Solve for a: 2a + 4 = 3a + 2 will give medal and fan 4 answer

OpenStudy (anonymous):

Subtract 2a from both sides 4=3a+2−2a Simplify 3a+2−2a to a+2 4=a+2 Subtract 2 from both sides 4−2=a Simplify 4−2 to 2 2=a Switch sides a=2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

a=2

OpenStudy (anonymous):

your welcome :)